OpenStudy (anonymous):

Find the area of the solid obtained by rotating the region bounded by the given curves about the specified line. y=2-.5x, y=0, x=1, x=2, about the x-axis

OpenStudy (campbell_st):

use \[\pi \int\limits_{1}^{2} y ^{2} dx\] or in your question \[\pi \int\limits_{1}^{2} (2 - 0.5x)^2 dx\] before integrating expand the quadratic and then integrate \[\pi \int\limits_{1}^{2} 4 - 2x + 1/4 x^2 dx\]

OpenStudy (anonymous):

thank you good sir or madam, thank you kindly.